OpenStudy (shelby1290):

Find an equation for the line parallel to 2x-3y+6=0, with the same y-intercept as y=7x-1

OpenStudy (anonymous):

Step 1: What is the y intercept of y = 7x -1 ?

OpenStudy (shelby1290):

-1 ?

OpenStudy (anonymous):

Yes... So the coordinates of that y-intercept is (0,-1). Does that make sense?

OpenStudy (shelby1290):

yes

OpenStudy (anonymous):

So our equation will have to go through that point. Now we just need to find the slope of our line. Our line will need to be parallel to 2x-3y+6=0 If it is gonna be parallel to that it needs to have the same slope as that. So....what is the slope of 2x-3y+6=0 ?

OpenStudy (shelby1290):

2 ?

OpenStudy (shelby1290):

2/3?

OpenStudy (anonymous):

No. To find the slope of that...put it in the form y = mx + b 2x - 3y + 6=0 2x - 3y =- 6 -3y = -2x - 6 .....now divide both sides by -3 y = 2/3 x + 2 So what's the slope ?

OpenStudy (anonymous):

yeah...made a small irrelevant mistake Slope is 2/3

OpenStudy (anonymous):

If we know a point on the line (0,1) and the slope... we can use the point slope form y - y1 = m(x - x1) x1 is 0 y1 is 1 m is 2/3 Plug all that in and you have your equation

OpenStudy (shelby1290):

how do i figure that out?

OpenStudy (shelby1290):

@BangkokGarrett ^^

OpenStudy (shelby1290):

the formula that I was taught is: y=m(x-p)+q

OpenStudy (anonymous):

Just plug in everything y - y1 = m(x - x1) y - 1 = 2/3(x - 0) y - 1 = 2/3 x y = 2/3 x + 1

OpenStudy (shelby1290):

ohhh okay i see

OpenStudy (anonymous):

ok...your formula is basically the same as mine...just arranged slightly differently